HassanAhmad:

\[x^{3})+y^{3})=(x+y)(x+wy)(x+w^{2}y)\]

HassanAhmad:

Need help!!!

XioGonz:

Do you mean (X^3+y^3)

HassanAhmad:

yesh i am new to this lol

XioGonz:

It's fine lol. You have to combine like terms

XioGonz:

In this case what are the like-terms that you can combine?

HassanAhmad:

man i can't even type the value of x here

XioGonz:

It's fine. Take your time

HassanAhmad:

do i have to combine \[x^{3}+y^{3}\] with \[y+x\]

HassanAhmad:

the question says prove that?

HassanAhmad:

well i will try it myself then lol

XioGonz:

My bad I was doing something.I Use this method. P() E x^2 M * D ÷ A + S -

HassanAhmad:

okay i will try thanks \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz My bad I was doing something.I Use this method. P() E x^2 M * D ÷ A + S - \(\color{#0cbb34}{\text{End of Quote}}\) okay i will try thanks for the help

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @HassanAhmad okay i will try thanks \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz My bad I was doing something.I Use this method. P() E x^2 M * D ÷ A + S - \(\color{#0cbb34}{\text{End of Quote}}\) okay i will try thanks for the help \(\color{#0cbb34}{\text{End of Quote}}\) Np teddy bear 😋

HassanAhmad:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @HassanAhmad okay i will try thanks \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz My bad I was doing something.I Use this method. P() E x^2 M * D ÷ A + S - \(\color{#0cbb34}{\text{End of Quote}}\) okay i will try thanks for the help \(\color{#0cbb34}{\text{End of Quote}}\) Np teddy bear 😋 \(\color{#0cbb34}{\text{End of Quote}}\) hahahahaha okay lol

jhonyy9:

(x^3+y^3)=(x+y)(x+wy)(x+w^2y) - @XioGonz so if you see the right part - my opinion - from there never not will get like result the x^3 +y^3 do you know it why ? bc. to get from there x^3 +y^3 so you need to have there somewhere an minus to get like result the rest of terms without x^3 and y^3 simplified (eliminated) hope i m understandably @supie

jhonyy9:

look pls how i ve learned much years ago x^3 +y^3 = (x + y)(x^2 -xy +y^2) = x^3 -x^2y +xy^2 +x^2y -xy^2 +y^3 = x^3 +y^3 hope helped understandably easy @supie

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 look pls how i ve learned much years ago x^3 +y^3 = (x + y)(x^2 -xy +y^2) = x^3 -x^2y +xy^2 +x^2y -xy^2 +y^3 = x^3 +y^3 hope helped understandably easy @supie \(\color{#0cbb34}{\text{End of Quote}}\) I agree.

Dixieland:

0.0

HassanAhmad:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) thank you i appreciate your help

jhonyy9:

np yw anytime

Dixieland:

this is confusing